# Tight PAC-Bayesian Risk Certificates for Contrastive Learning

### v2

@@ Results
Table 2 demonstrates that our proposed risk certificates for the SimCLR loss are non-vacuous and significantly outperform existing risk certificates on CIFAR-10, closely aligning with the corresponding test losses. Interestingly, theorem 4 yields tighter certificates for $\tau\leq 0.5$, while theorem 3 is more effective for $\tau>0.5$. We also observe that Catoni’s bound is significantly tighter than the PAC-Bayes-kl bound, which is consistent with its known advantage when $\operatorname{KL}/n$ is large [47]. Unsurprisingly, the classic PAC-Bayes bound is looser than both of these bounds. Additionally, Table 3 shows that theorem 5 improves upon the bound from Bao et al., which results in exponential growth when $\tau\leq 0.5$ [4]. Moreover, models trained using PAC-Bayes by Backprop achieve competitive top1 accuracy. Table 2 further illustrates that our risk certificates for contrastive zero-one risk are notably tight, surpassing existing certificates. Additionally, we observe that theorem 7 consistently outperforms theorem 6. Overall, our risk certificates are competitive, even for low temperatures—a known challenge. In PAC-Bayes learning, we observe that a randomly initialized prior leads to poor convergence and sub-optimal risk certificates, while an informed prior results in better performance and tighter certificates. However, PAC-Bayes learning in models with a large number of parameters remains challenging and warrants further investigation (most studies focus on 2 or 3 hidden layers, while our numerical results are on a 7-layer network).
@@ 5 Discussion and Future Work
We have presented novel PAC-Bayesian risk certificates tailored for the SimCLR framework. Our experiments on CIFAR-10 show that our bounds yield non-vacuous risk certificates and significantly outperform previous ones.
Bounding techniques. theorem 4 and theorem 7 rely on concentration bounds to apply the PAC-Bayes bound-kl in an i.i.d. setting. Although the PAC-Bayes-kl bound was selected for its tightness, any bound that respects our proof’s assumptions could be applicable. Moreover, theorem 3 and theorem 6 were obtained using McDiarmid’s inequality, integrated into McAllester’s PAC-Bayes bound. Since this variant is known to be less tight than the kl bound or Catoni’s bound, it would be interesting to explore whether McDiarmid’s inequality can be incorporated into these tighter bounds.
Extension to other (non-i.i.d.) losses While this paper primarily focuses on contrastive learning using the SimCLR framework, the approaches we propose to address the non-i.i.d. characteristics of the SimCLR loss can be readily applied to other loss presenting similar dependence, such as ranking losses [8], Barlow Twins [46] or VICReg [5]. While our non-i.i.d. McAllester PAC-Bayes bound (theorem 3) requires only a bounded difference assumption, the non-i.i.d. PAC-Bayes-kl bound (theorem 4), while more opaque, requires only a Hoeffding’s assumption on the dependent terms.
Impact of temperature scaling and projection head Temperature scaling in the SimCLR loss remains challenging as the scaling constant loosens the PAC-Bayes bounds, suggesting the need for more adapted PAC-Bayes bounds for this type of loss. Regarding the bound on the downstream classification loss, our approach better handles smaller temperatures than previous bounds, though it still struggles to perfectly align with downstream classification losses at low temperatures. Regarding projection head, we empirically observe that classification loss is lower when features are used without a projection head compared to with one, yet the theoretical role of the projection head cannot fully be understood with our downstream classification bound. In this work, we proposed an approach to integrate a simple projection head into our bound, and it would be interesting to link this with [22] that suggests a fixed low-rank diagonal projector might suffice instead of a trainable projection head.
PAC-Bayes learning Our models trained with PAC-Bayes by Backprop achieve accuracy competitive with [7] despite using a much smaller model (7-layer CNN vs. 50-layer CNN) and training for fewer epochs (100 vs. 500). Although this paper focuses on deriving better risk certificates rather than improving the PAC-Bayes learning algorithm, there is a need to extend the PAC-Bayes paradigm to large-scale neural networks, such as ResNet50.
@@ B.3 Experiments on MNIST
In this section, we present the results of the experiments on MNIST, as detailed in table 4, table 5, and table 6.
Risk Certificate SimCLR Loss
@@
Risk Certificate Contrastive 0-1 Risk

### v4

@@ Results
Table 2 demonstrates that our proposed risk certificates for the SimCLR loss are non-vacuous and significantly outperform existing risk certificates on CIFAR-10, closely aligning with the corresponding test losses. Interestingly, theorem˜4 yields tighter certificates for $\tau\leq 0.5$, while theorem˜3 is more effective for $\tau>0.5$. We also observe that Catoni’s bound is significantly tighter than the PAC-Bayes-kl bound, which is consistent with its known advantage when $\operatorname{KL}/n$ is large [58]. Unsurprisingly, the classic PAC-Bayes bound is looser than both of these bounds. Additionally, Table 3 shows that theorem˜5 improves upon the bound from Bao et al., which results in exponential growth when $\tau\leq 0.5$ [3]. Moreover, models trained using PAC-Bayes by Backprop achieve competitive top1 accuracy. Table 2 further illustrates that our risk certificates for contrastive zero-one risk are notably tight, surpassing existing certificates. Additionally, we observe that theorem˜7 consistently outperforms theorem˜6. Overall, our risk certificates are competitive, even for low temperatures. PAC-Bayes learning in models with a large number of parameters remains challenging and warrants further investigation (most studies focus on 2 or 3 hidden layers).
@@ 5 Discussion and Future Work
We have presented novel PAC-Bayesian risk certificates tailored for the SimCLR framework. Our experiments on CIFAR-10 and MNIST show that our bounds yield non-vacuous risk certificates and significantly outperform previous ones.
Bounding techniques
theorem˜4 and theorem˜7 rely on concentration bounds to apply the PAC-Bayes bound-kl in an i.i.d. setting. Although the PAC-Bayes-kl bound was selected for its tightness, any bound that respects our proof’s assumptions could be applicable. Moreover, theorem˜3 and theorem˜6 were obtained using McDiarmid’s inequality, integrated into McAllester’s PAC-Bayes bound. Since this variant is known to be less tight than the kl bound or Catoni’s bound, it would be interesting to explore whether McDiarmid’s inequality can be incorporated into these tighter bounds.
Extension to other (non-i.i.d.) losses
While this paper primarily focuses on contrastive learning using the SimCLR framework, the approaches we propose to address the non-i.i.d. characteristics of the SimCLR loss can be readily applied to other loss presenting similar dependence, such as ranking losses [8], Barlow Twins [56] or VICReg [5]. While our non-i.i.d. McAllester PAC-Bayes bound (theorem˜3) requires only a bounded difference assumption, the non-i.i.d. PAC-Bayes-kl bound (theorem˜4), while more opaque, requires only a Hoeffding’s assumption on the dependent terms.
Impact of temperature scaling and projection head
Temperature scaling in the SimCLR loss remains challenging as the scaling constant loosens the PAC-Bayes bounds, suggesting the need for more adapted PAC-Bayes bounds for this type of loss. Regarding the bound on the downstream classification loss, our approach better handles smaller temperatures than previous bounds, though it still struggles to perfectly align with downstream classification losses at low temperatures.
Regarding projection head, we empirically observe that classification loss is lower when features are used without a projection head compared to with one, yet the theoretical role of the projection head cannot fully be understood with our downstream classification bound. In this work, we proposed an approach to integrate a simple projection head into our bound, and it would be interesting to link this with [26] that suggests a fixed low-rank diagonal projector might suffice instead of a trainable projection head.
PAC-Bayes learning
Our models trained with PAC-Bayes by Backprop achieve accuracy competitive with [7] despite using a much smaller model (7-layer CNN vs. 50-layer CNN) and training for fewer epochs (100 vs. 500). Although this paper focuses on deriving better risk certificates rather than improving the PAC-Bayes learning algorithm, there is a need to extend the PAC-Bayes paradigm to large-scale neural networks, such as ResNet50.
Acknowledgments
We would like to express our gratitude to Maximilian Fleissner for providing Lemma 3.14 and for his careful reading of the original manuscript. We also sincerely thank the SIMODS reviewers for their valuable comments and constructive feedback, which greatly improved the quality of this work.
@@ B.3 Experiments on MNIST
In this section, we present the results of the experiments on MNIST, as detailed in table˜4, table˜5, and table˜6.
Risk Certificate
SimCLR Loss
@@
Risk Certificate
Contrastive 0-1 Risk
@@ Appendix C KL Divergence
Appendix D Detailed Proof of Lemma 3.20
Below, we provide a detailed proof of lemma˜11. Specifically, we aim to show that
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\sigma+\tau L(f)+\tau\Delta+\alpha\,$,
where $\alpha=\log(C)+\min\left\{0,\log\left(C\pi^{*}\cosh^{2}(1)\right)-\tau\Delta\right\}$.
Proof.
On one hand, using two applications of (23), we obtain an upper bound on the term $\operatorname{LSE}(\mathbf{u})$ with $\mathbf{u}:=\left\{f(x)^{\top}f(x^{\prime})\right\}_{x^{\prime}\in X}$:
$\operatorname{LSE}(\mathbf{u})=\operatorname{LSE}\left(\tau\cdot\frac{\mathbf{u}}{\tau}\right)\leq\tau\max_{i}z_{i}+\log(K)\leq\tau\operatorname{LSE}(\mathbf{z})+\log(K),$ (26)
where the first inequality uses $t=\tau$ and the second one uses $t=1$. This yields:
$\operatorname{LSE}(\mathbf{z})\geq\frac{1}{\tau}\left(\operatorname{LSE}(\mathbf{u})-\log(K)\right).$ (27)
We obtain:
$L(f)$ $\overset{(a)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{(x_{i}^{\prime},y_{i}^{\prime})\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z})\right]$
$\overset{(b)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\frac{1}{\tau}\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{(x_{i}^{\prime},y_{i}^{\prime})\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{u})\right]-\frac{\log(K)}{\tau}$
$\overset{(c)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\frac{1}{\tau}\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{y_{i}^{\prime}\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{u}^{\mu})\right]-\frac{\log(K)}{\tau},$
where: (a) follows from the sequence of inequalities in eq.˜11 to eq.˜14; (b) follows by applying eq.˜27; (c) follows from an application of Jensen’s inequality with $\mathbf{u}^{\mu}:=\left\{f(x)^{\top}\mu_{y^{\prime}}\right\}_{y^{\prime}\in Y}$.
Then, we derive a lower bound on $\mathbb{E}_{Y}\left[\operatorname{LSE}(\mathbf{u}^{\mu})\right]$:
$\mathbb{E}_{Y}\left[\operatorname{LSE}(\mathbf{u}^{\mu})\right]$ $\overset{(a)}{\geq}-\mathbb{E}_{Y}\left[\operatorname{LSE}(-\mathbf{u}^{\mu})\right]+2\log(K)$
$\overset{(b)}{\geq}-\operatorname{LSE}\left(\left\{-f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)+\log(K)-\log(\pi^{*})$
$\overset{(c)}{\geq}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)-2\log\left(C\cosh(1)\right)+\log(K)-\log(\pi^{*})$
$\overset{(d)}{=}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right)-\log\left(\frac{C^{2}\pi^{*}}{K}\cosh^{2}(1)\right),$
where (a) uses the left-hand side of lemma˜10 for $N=K$ and $L=1$, (b) is a direct application of eq.˜17 to eq.˜15, using $\mathbf{u}^{\mu}$ instead of $\mathbf{z}^{\mu}$ (they differ only by a scaling factor $\tau$), (c) applies the right-hand side of lemma˜10 for $N=C$ and $L=1$, (d) follows by simplifying the previous line.
Combining this with the earlier bound, we obtain:
$L(f)$ $\geq{-}\frac{\sigma}{\tau}{-}\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]{+}\frac{1}{\tau}\operatorname{LSE}\left(\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}\right){-}\frac{1}{\tau}\log\left(\frac{C^{2}\pi^{*}}{K}\cosh^{2}(1)\right){-}\frac{\log(K)}{\tau}$
$=-\frac{\sigma}{\tau}+\frac{1}{\tau}L_{\mathrm{CE}}(f,W^{\mu})-\frac{1}{\tau}\log\left(C^{2}\pi^{*}\cosh^{2}(1)\right).$
Thus, we obtain the following bound on the mean classifier:
$L_{\mathrm{CE}}(f,W^{\mu})\leq\sigma+\tau L(f)+\tau\Delta+\log(C)+\log\left(C\pi^{*}\cosh^{2}(1)\right)-\tau\Delta.$ (28)
On the other hand, applying (23) twice, we obtain an upper bound on the term $\operatorname{LSE}(\mathbf{u}^{c})$, where $\mathbf{u}^{c}:=\left\{f(x)^{\top}\mu_{c}\right\}_{c\in\mathcal{C}}$:
$\operatorname{LSE}(\mathbf{u}^{c})=\operatorname{LSE}\left(\tau\cdot\frac{\mathbf{u}^{c}}{\tau}\right)\leq\tau\max_{i}z_{i}^{c}+\log C\leq\tau\operatorname{LSE}(\mathbf{z}^{c})+\log C,$ (29)
where we define $\operatorname{LSE}(\mathbf{z}^{c})$ with $\mathbf{z}^{c}:=\left\{\frac{f(x)^{\top}\mu_{c}}{\tau}\right\}_{c\in\mathcal{C}}$. This yields
$\tau\operatorname{LSE}(\mathbf{z}^{c})\geq\operatorname{LSE}(\mathbf{u}^{c})-\log C.$ (30)
We obtain:
$L(f)$ $\overset{(a)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\mathbb{E}_{x}\mathbb{E}_{\begin{subarray}{c}\{(x_{i}^{\prime},y_{i}^{\prime})\}_{i=1}^{K}\end{subarray}}\left[\operatorname{LSE}(\mathbf{z}^{\mu})\right]$
$\overset{(b)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\operatorname{LSE}(\mathbf{z}^{c})-\Delta$
$\overset{(c)}{\geq}-\frac{\sigma}{\tau}-\mathbb{E}_{(x,y)}\left[\frac{f(x)^{\top}\mu_{y}}{\tau}\right]+\frac{1}{\tau}\operatorname{LSE}(\mathbf{u}^{c})-\frac{\log C}{\tau}-\Delta$
where: (a) follows from the sequence of inequalities from eq.˜11 to eq.˜15; (b) follows from applying eq.˜16 to eq.˜22; (c) follows by applying eq.˜30. We obtain
$L_{\mathrm{CE}}(f,W^{\mu})\leq\sigma+\tau L(f)+\tau\Delta+\log C.$ (31)
Finally, taking the minimum over the two refined bounds eq.˜28 and eq.˜31, and then over all linear classifiers completes the proof. ∎
Appendix E Extensions
Throughout the paper, we have noted several directions for generalizing our results to better align with the original SimCLR framework. In this section, we elaborate on these remarks and provide explicit derivations to support each extension.
Remark 2.1
We now elaborate on the statement made in remark˜1. Specifically, we detail the formulation of the original SimCLR loss introduced in [7], applied over a dataset $S\sim\mathcal{S}^{n}$:
$\widehat{L}_{S}(f)=\frac{1}{n}\sum_{i=1}^{n}\frac{\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},X_{i}^{-})+\ell_{\operatorname{cont}}(x_{i}^{+},x_{i},X_{i}^{-})}{2}=\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},x_{i}^{+},X_{i}^{-}),$ (32)
where we recall that the set of negatives is given by $X_{i}^{-}=\bigcup_{j\neq i}\{x_{j},x_{j}^{+}\}$.
Remark 3.3
We now provide details on how the proof of lemma˜1 adapts to the loss defined in eq.˜32, see remark˜2. The extension requires only minor modifications:
  1. i)
    The set $X_{i}$ is replaced with the full set of 2(m1)2(m-1) negative samples, denoted $X_{i}^{-}$.
  2. ii)
    The set $N_{i}$ is replaced with $N_{i}^{-}=X_{i}^{-}\!\setminus\!\{x_{i},x_{i}^{+}\}$.
  3. iii)
    The quantities $\kappa$, $a$, and $b$ are redefined appropriately.
As a consequence, we obtain the following upper bound:
$\left|\ell_{\operatorname{cont}}(x_{i},x_{i}^{+},X_{i}^{-})-\ell_{\operatorname{cont}}(x_{i}^{\prime},x_{i}^{\prime+},X_{i}^{-})\right|\leq\frac{4}{\tau}.$
Let $\widetilde{X}_{j}^{-}$ denote a perturbed version of the negative set $X_{j}^{-}$. Then:
$\left|\ell_{\operatorname{cont}}(x_{j},x_{j}^{+},X_{j}^{-})-\ell_{\operatorname{cont}}(x_{j},x_{j}^{+},\widetilde{X}_{j}^{-})\right|=\left|\log\frac{\kappa+\operatorname{sim}(x_{j},x_{i})+\operatorname{sim}(x_{j},x_{i}^{+})}{\kappa+\operatorname{sim}(x_{j},x_{i}^{\prime})+\operatorname{sim}(x_{j},{x_{i}^{\prime}}^{+})}\right|\leq\log\frac{\kappa+a}{\kappa+b},$
where $\kappa=\operatorname{sim}(x_{j},x_{j}^{+})+\sum_{x^{\prime}\in N_{i}^{-}}\operatorname{sim}(x_{j},x^{\prime})$, $a=2e^{1/\tau}$ and $b=2e^{-1/\tau}$. Since $\kappa$ includes 2m32m-3 terms, it is lower-bounded as $\kappa\geq(2m-3)e^{-1/\tau}$. Thus, we can further bound:
$\log\frac{\kappa+a}{\kappa+b}\leq\log\frac{(2m-3)e^{-1/\tau}+2e^{1/\tau}}{(2m-3)e^{-1/\tau}+2e^{-1/\tau}}=\log\frac{(2m-3)+2e^{2/\tau}}{2m-1}.$
Arguing similarly to the proof of lemma˜1 we conclude that
$|\delta_{i}(\bar{x}_{i})|\leq\frac{4}{\tau}\quad\text{and}\quad|\delta_{j}(\bar{x}_{i})|\leq\log\frac{(2m-3)+2e^{2/\tau}}{2m-1}.$
Combining these yields $\frac{4}{\tau}+(m-1)\log\frac{(2m-3)+2e^{2/\tau}}{2m-1}$.
Remark 3.9
We now detail how the proof of theorem˜4 extends to the loss defined in eq.˜32; see also remark˜3. In this extension, the set $X$ is replaced by $X^{-}$, and we therefore require a concentration bound on the quantity
$S(x,X^{-}):=\sum\limits_{x^{\prime}\in X^{-}}\operatorname{sim}(x,x^{\prime})=\sum\limits_{x^{\prime}\in X}\left(\operatorname{sim}(x,x^{\prime})+\operatorname{sim}(x,{x^{\prime}}^{+})\right).$
This expression shows that $S(x,X^{-})$ is a sum of m1m-1 independent and bounded random variables, each lying within the interval $2e^{-1/\tau},2e^{1/\tau}]$. We apply Hoeffding’s inequality with range parameter $c=2\left(e^{1/\tau}-e^{-1/\tau}\right)$, and obtain the same concentration bound as in lemma˜5, with the only difference being that the term $\varepsilon$ is now given by $\varepsilon=2\left(e^{1/\tau}-e^{-1/\tau}\right)\sqrt{\frac{m-1}{2}\log\frac{1}{\delta}},$ which is multiplied by a factor of 2 compared to the previous result.
Remark 3.16
As discussed in remark˜4, theorem˜5 can be extended to include a simple projection head. Specifically, we first apply the theorem to the projected features $f_{1}$, which yields
$\min_{W^{(1)}\in\mathbb{R}^{C\times k}}L_{\mathrm{CE}}(f_{1},W^{(1)})\leq\min\left\{\beta(f_{1},\sigma),~\tau\beta(f_{1},\sigma)+\alpha\right\},$
[…]
$\tilde{W}=\begin{bmatrix}W^{(1)}\\ 0_{(d-k)\times C}\end{bmatrix}$,
[…]
$\min_{W\in\mathbb{R}^{C\times d}}L_{\mathrm{CE}}(f,W)\leq\min\left\{\beta(f_{1},\sigma),~\tau\beta(f_{1},\sigma)+\alpha\right\}.$
Remark 3.17
As mentioned in remark˜5, the proof of theorem˜5 can be directly extended to the loss defined in eq.˜32 by replacing $X$ with $X^{-}$ and setting K=2(m1)K=2(m-1) throughout the derivation. Note that the proof remains unchanged, as it does not rely on any independence assumptions regarding the negative samples.